[Hallc_running] [New Logentry] RC daily update

billlee at jlab.org billlee at jlab.org
Mon Sep 17 10:00:05 EDT 2018

Logentry Text:
>From the MCC meeting this morning:
--> Accelerator components start to lock-up starting around lunch time today. SL could restore RF today. Aperture test is planed for tomorrow (Sep/18).
--> 65 uA tune (CW)  beam into Hall C by Thursday Swing shift.

Hall C update:
--> Isopropanol Refrigerator door was left open by me (really sorry about that),  when performing the checklist on Friday. Temperature raised and  ice formed inside of the refrigerator. Thanks to Howard that the refrigerator is defrosted and working again.

--> All systems remain shut. Experts are expected to come in and start to turn things on.
--> Howard restored DC normal gas flow this morning and DC will be ready for operation by Tuesday noon. A cosmic ray test is planned Tuesday afternoon

--> The HMS and SHMS motors were installed and  electrically connected, a test is planed (expected) this morning.
--> Hall lock down is expected by the end of the day-shift next Thursday (4pm, Sep/20th). 
--> Shifts remain canceled until 8am, Sep/20th (Thursday).
--> We will have a RC meeting today.

Planned work:
--> Steve and Joe will test the newly install motors.
--> Beam permit check list will be performed early this week.
--> A cosmic ray test is planned for next Tuesday (Sep/18) afternoon.
--> Target ladder cool down early this week.


This is a plain text email for clients that cannot display HTML.  The full logentry can be found online at https://logbooks.jlab.org/entry/3594664
-------------- next part --------------
An HTML attachment was scrubbed...
URL: <https://mailman.jlab.org/pipermail/hallc_running/attachments/20180917/72f1219e/attachment.html>

More information about the Hallc_running mailing list